OpenStudy (el_arrow):

whats the derivative of 5(1+cos(x))?

OpenStudy (el_arrow):

is it -5sinx?

OpenStudy (anonymous):

yes

OpenStudy (bibby):

d/dx cf(x)=cf'(x) \(\dfrac{d}{dx}5(1+cos(x))=5\dfrac{d}{dx}(1+cos(x))=5(\dfrac{d}{dx}1+\dfrac{d}{dx}cos(x)\))

OpenStudy (bibby):

5(0+-sin(x))=-5sinx, yea

OpenStudy (el_arrow):

okay thank you